The owner of a fast-food shop claims more than 45% of the household within 5-kilometer radius of the fast-food shop have at least one member who eats at the fast-food shop at least once a week. In a sample of 250 households in the area, an investigator found that members of 180 households did so. Do these data provide sufficient evidence to support the fast-food shop owner’s claim? Use 90% level of confidence.

1
Expert's answer
2022-06-01T02:42:05-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le0.45"

"H_a:p>0.45"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.10

," and the critical value for a right-tailed test is "z_c =1.2816."

The rejection region for this right-tailed test is "R = \{z: z > 1.2816\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{180/250-0.45}{\sqrt{\dfrac{0.45(1-0.45)}{250}}}=8.5812"

Since it is observed that "z = 8.5812>1.2816= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>8.5812)=0," and since "p=0<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than 0.45, at the "\alpha = 0.10" significance level.